# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ D.3.4 The Myth of Sisyphus, Camus
Figure 20: The Myth of Sisyphus, Camus [13]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.5 Alice’s Adventures in Wonderland, Carroll
@@
Figure 22: Alice’s Adventures in Wonderland, Carroll [19]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.6 The Beautiful Struggle, Coates
@@
D.3.7 We Were Eight Years in Power, Coates
@@
Figure 25: We Were Eight Years in Power, Coates [23]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.8 The Water Dancer, Coates
@@
Figure 27: The Water Dancer, Coates [24]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.9 Dungeons and Dragons and Philosophy, Cogburn
@@
Figure 29: Dungeons and Dragons and Philosophy, Cogburn [25]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.10 Drown, D´ıaz
Figure 31: Drown, D´ıaz [38]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.11 The Brief Wondrous Life of Oscar Wao, D´ıaz
@@
Figure 33: The Brief Wondrous Life of Oscar Wao, D´ıaz [39]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.12 This Is How You Lose Her, D´ıaz
Figure 35: This Is How You Lose Her, D´ıaz [40]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.13 Down and Out in the Magic Kingdom, Doctorow
@@
Figure 37: Down and Out in the Magic Kingdom, Doctorow [36]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.14 The World’s Wife, Duffy
@@
Figure 39: The World’s Wife, Duffy [37]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.15 A Visit from the Goon Squad, Egan
@@ A Visit from the Goon Squad
Figure 41: A Visit from the Goon Squad, Egan [41]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.16 The President’s Vampire, Farnsworth
@@
Figure 43: The President’s Vampire, Farnsworth [43]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.17 The Great Gatsby, Fitzgerald
@@
Figure 45: The Great Gatsby, Fitzgerald [46]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.18 Blink, Gladwell
Figure 47: Blink, Gladwell [49]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.19 Dead Ringer, Golden
@@
Figure 49: Dead Ringer, Golden [50]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.20 Ararat, Golden
Figure 51: Ararat, Golden [51]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.21 The Confessions of Max Tivoli, Greer
@@
Figure 53: The Confessions of Max Tivoli, Greer [55]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.22 The Fugitive, Grisham
Figure 55: The Fugitive, Grisham [56]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.23 The Curious Incident of the Dog in the Night-Time, Haddon
@@
Figure 57: The Curious Incident of the Dog in the Night-Time, Haddon [57]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.24 Catch-22, Heller
Figure 59: Catch-22, Heller [61]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
Figure 61: M. Butterfly, Hwang [65]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.26 All the Onions, Jacobs
Figure 63: All the Onions, Jacobs [67]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.27 Ulysses, Joyce
Figure 65: Ulysses, Joyce [68]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.28 Sandman Slim, Kadrey
@@
Figure 67: Sandman Slim, Kadrey [69]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.

### v4

@@ D.3.4 The Myth of Sisyphus, sisyphus
Figure 20: The Myth of Sisyphus, sisyphus: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.5 Alice’s Adventures in Wonderland, alice
@@
Figure 22: Alice’s Adventures in Wonderland, alice: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.6 The Beautiful Struggle, struggle
@@
D.3.7 We Were Eight Years in Power, wewere
@@
Figure 25: We Were Eight Years in Power, wewere: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.8 The Water Dancer, waterdancer
@@
Figure 27: The Water Dancer, waterdancer: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.9 Dungeons and Dragons and Philosophy, dungeons
@@
Figure 29: Dungeons and Dragons and Philosophy, dungeons: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.10 Drown, drown
Figure 31: Drown, drown: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.11 The Brief Wondrous Life of Oscar Wao, oscarwao
@@
Figure 33: The Brief Wondrous Life of Oscar Wao, oscarwao: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.12 This Is How You Lose Her, loseher
Figure 35: This Is How You Lose Her, loseher: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.13 Down and Out in the Magic Kingdom, downandout
@@
Figure 37: Down and Out in the Magic Kingdom, downandout: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.14 The World’s Wife, worldswife
@@
Figure 39: The World’s Wife, worldswife: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.15 A Visit from the Goon Squad, goonsquad
@@ A Visit from the Goon Squad
Figure 41: A Visit from the Goon Squad, goonsquad: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.16 The President’s Vampire, vampire
@@
Figure 43: The President’s Vampire, vampire: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.17 The Great Gatsby, gatsby
@@
Figure 45: The Great Gatsby, gatsby: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.18 Blink, blink
Figure 47: Blink, blink: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.19 Dead Ringer, deadringer
@@
Figure 49: Dead Ringer, deadringer: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.20 Ararat, ararat
Figure 51: Ararat, ararat: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.21 The Confessions of Max Tivoli, maxtivoli
@@
Figure 53: The Confessions of Max Tivoli, maxtivoli: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.22 The Fugitive, fugitive
Figure 55: The Fugitive, fugitive: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.23 The Curious Incident of the Dog in the Night-Time, curious
@@
Figure 57: The Curious Incident of the Dog in the Night-Time, curious: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.24 Catch-22, catch22
Figure 59: Catch-22, catch22: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
Figure 61: M. Butterfly, mbutterfly: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.26 All the Onions, onions
Figure 63: All the Onions, onions: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.27 Ulysses, ulysses
Figure 65: Ulysses, ulysses: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.28 Sandman Slim, sandmanslim
@@
Figure 67: Sandman Slim, sandmanslim: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.